A report says that the average energy content of candy bars is between 220 and 314 calories. What is the margin of error in this report?

Select one or more:

a. ME = 104

b. ME = 47

c. ME = 96

d. ME = 64.5
Answer #1

Solution :

\bar x  = (Lower bound + Upper bound ) / 2

Sample mean = \bar x = (220+314) / 2=267

Margin of error = E = Upper bound - \bar x   = 314-267=47

b. ME = 47